# The Super Weight in Large Language Models

### v1

@@ Abstract
Recent works have shown a surprising result: a small fraction of Large Language Model (LLM) parameter outliers are disproportionately important to the quality of the model. LLMs contain billions of parameters, so these small fractions, such as 0.01%, translate to hundreds of thousands of parameters. In this work, we present an even more surprising finding: Pruning as few as a single parameter can destroy an LLM’s ability to generate text – increasing perplexity by 3 orders of magnitude and reducing zero-shot accuracy to guessing. We propose a data-free method for identifying such parameters, termed super weights, using a single forward pass through the model. We additionally find that these super weights induce correspondingly rare and large activation outliers, termed super activations. When preserved with high precision, super activations can improve simple round-to-nearest quantization to become competitive with state-of-the-art methods. For weight quantization, we similarly find that by preserving the super weight and clipping other weight outliers, round-to-nearest quantization can scale to much larger block sizes than previously considered. To facilitate further research into super weights, we provide an index of super weight coordinates for common, openly available LLMs [^1] Code is available in https://github.com/mengxiayu/LLMSuperWeight..
@@ 1 Introduction
Figure 1: Super Weight Phenemenon. We discover that pruning a single, special scalar, which we call the super weight, can completely destroy a Large Language Model’s ability to generate text. On the left, the original Llama-7B, which contains a super weight, produces a reasonable completion. On the right, after pruning the super weight, Llama-7B generates complete gibberish. As we show below, this qualitative observation has quantitative impact too: zero-shot accuracy drops to guessing and perplexity increases by orders of magnitude.
@@ 3.1 Identification of Super Weights
Suppose that we have a down_proj weight matrix $\mathbf{W}\in\mathbb{R}^{D\times H}$, where $D$ is the dimension of the activation feature and $H$ is the intermediate hidden dimension. Let $\mathbf{X}\in\mathbb{R}^{L\times H}$ be the input matrix, where $L$ is the sequence length. $\mathbf{Y}=\mathbf{X}\mathbf{W}^{T}$, where $\mathbf{Y}_{ij}=\sum_{k=1}^{d}\mathbf{X}_{ik}\mathbf{W}_{jk}$. Suppose $\mathbf{Y}{ij}$ is a super activation. If $\mathbf{X}{ik}$ and $\mathbf{W}{jk}$ are both outliers that are much larger than other values, $\mathbf{Y}{ij}$ will be dominated by their product. That is, $\mathbf{Y}{ij}$approx $\mathbf{X}{ik}$ $\mathbf{W}{jk}$. In this case, $j$ and $k$ are determined by $\mathbf{X}{ik}$ and $\mathbf{Y}{ij}$. Therefore, we start by plotting extreme outliers in the input and output activations of mlp.down_proj. Then, we determine the layer and coordinates of the super weight, as illustrated in Figure  4. Once we have detected one super weight, we remove it from the model and repeat the above process, until the magnitudes of large maximum activations are greatly suppressed.
@@
Figure 2: How Super Weights behave. I: Super weights are often found in an early layer’s down projection, indicated with a blue-purple box. The super weight immediately creates an incredibly large-magnitude super activation. II: Super activations are propagated through skip connections, indicated with blue-purple lines. III: This has a net effect of suppressing stopword likelihoods in the final logits. Removing the super weight causes stopword likelihood skyrocket, indicated with the blue-purple stacked bars. See Appendix A.3.
@@ 3.2 Mechanisms of super weights
We conduct an ablation experiment with three conditions: (1) the original model, (2) remove the super weight (Prune SW), i.e., setting the weight scalar as zero, (3) remove the super weight and restore the super activation at the layer where it first appears (Prune SW,+SA). The third condition allows us to isolate the impact of super weights on super activations only.
@@
We aim to illustrate how variations in the magnitude of super weights impact the model’s quality, especially, how does increasing the magnitude affect model quallity. We multiply the super weights by a scaling factor ranging from 0.0 to 3.0. Results in Figure 6 show that amplifying super weights can improve model accuracy, to some extent. See full versions of these plots, for more models and all datasets, in Appendix A.1.
@@ A.1 Super Weight Sensitivity
In this section, we show the full set of results on zero-shot downstream datasets. Results are shown in Table 5 for FP16 models. From the table, we can see that some datasets are more sensitive to super weight (SW) amplification. For example, Winogrande and Lambada show consistent improvements across models when amplifying SW, while PIQA shows same or slightly lower accuracy. We also evaluate the 4-bit quantized Llama-7B with amplified SW in Table 6. We witness similar minor improvements when SW is amplified.

### v2

@@ Abstract
Recent works have shown a surprising result: a small fraction of Large Language Model (LLM) parameter outliers are disproportionately important to the quality of the model. LLMs contain billions of parameters, so these small fractions, such as 0.01%, translate into hundreds of thousands of parameters. In this work, we present an even more surprising finding: Pruning as few as a single parameter can destroy an LLM’s ability to generate text – increasing perplexity by 3 orders of magnitude and reducing zero-shot accuracy to guessing. We propose a data-free method for identifying such parameters, termed super weights, using a single forward pass through the model. We additionally find that these super weights induce correspondingly rare and large activation outliers, termed super activations. When preserved with high precision, super activations can improve simple round-to-nearest quantization to become competitive with state-of-the-art methods. For weight quantization, we similarly find that by preserving the super weight while clipping other weight outliers, round-to-nearest quantization can scale to much larger block sizes than previously considered. To facilitate further research into super weights, we provide an index of super weight coordinates for common, openly available LLMs.
@@ 1 Introduction
Figure 1: Super Weight Phenomenon. We discover that pruning a single, special scalar, which we call the super weight, can completely destroy a Large Language Model’s ability to generate text. On the left, the original Llama-7B, which contains a super weight, produces a reasonable completion. On the right, after pruning the super weight, Llama-7B generates complete gibberish. As we show below, this qualitative observation has quantitative impact too: zero-shot accuracy drops to guessing and perplexity increases by orders of magnitude.
@@ 3.1 Identification of Super Weights
Suppose that we have a down_proj weight matrix $\mathbf{W}\in\mathbb{R}^{D\times H}$, where $D$ is the dimension of the activation feature and $H$ is the intermediate hidden dimension. Let $\mathbf{X}\in\mathbb{R}^{L\times H}$ be the input matrix, where $L$ is the sequence length. $\mathbf{Y}=\mathbf{X}\mathbf{W}^{T}$, where $\mathbf{Y}_{ij}=\sum_{k=1}^{d}\mathbf{X}_{ik}\mathbf{W}_{jk}$. Suppose $\mathbf{Y}_{ij}$ is a super activation. If $\mathbf{X}_{ik}$ and $\mathbf{W}_{jk}$ are both outliers that are much larger than other values, $\mathbf{Y}{ij}$ will be dominated by their product. That is, $\mathbf{Y}_{ij}$approx $\mathbf{X}_{ik}$ $\mathbf{W}_{jk}$. In this case, $j$ and $k$ are determined by $\mathbf{X}_{ik}$ and $\mathbf{Y}_{ij}$. Therefore, we start by plotting extreme outliers in the input and output activations of mlp.down_proj. Then, we determine the layer and coordinates of the super weight, as illustrated in Figure  4. Once we have detected one super weight, we remove it from the model and repeat the above process, until the magnitudes of large maximum activations are greatly suppressed.
@@
Figure 2: How Super Weights behave. I: Super weights are often found in an early layer’s down projection, indicated with a blue-purple box. The super weight immediately creates an incredibly large-magnitude super activation. II: Super activations are propagated through skip connections, indicated with blue-purple lines. III: This has a net effect of suppressing stopword likelihoods in the final logits. Removing the super weight causes stopword likelihood to increase, indicated with the blue-purple stacked bars. See Appendix A.3.
@@ 3.2 Mechanisms of super weights
We conduct an ablation experiment under three conditions: (1) the original model, (2) remove the super weight (Prune SW), i.e., setting the weight scalar as zero, (3) remove the super weight and restore the super activation at the layer where it first appears (Prune SW,+SA). The third condition allows us to isolate the impact of super weights on super activations only.
@@
We aim to illustrate how variations in the magnitude of super weights impact the model’s quality, especially, how does increasing the magnitude affect model quality. We multiply the super weights by a scaling factor ranging from 0.0 to 3.0. Results in Figure 6 show that amplifying super weights can improve model accuracy, to some extent. See full versions of these plots, for more models and all datasets, in Appendix A.1.
@@ A.1 Super Weight Sensitivity
In this section, we show the full set of results on zero-shot downstream datasets. Results are shown in Table 5 for FP16 models. From the table, we can see that some datasets are more sensitive to super weight (SW) amplification. For example, Winogrande and Lambada show consistent improvements across models when amplifying SW, while PIQA shows same or slightly lower accuracy. We also evaluate the 4-bit quantized Llama-7B with amplified SW in Table 6. We observe similar minor improvements when SW is amplified.